# SciNUP: Natural Language User Interest Profiles for Scientific Literature Recommendation

### v1

@@ 4 The SciNUP Dataset
We present SciNUP (Scientific Natural Language User Profiles), a synthetic dataset designed for evaluating natural language-based recommendation in the scholarly domain. We envision a scenario where a researcher is looking for article recommendations relevant to their interests. Our core idea to simulate this setting by leveraging the individual’s publication history: their past publications are used to create NL user interest profiles, while cited references in their future works serve as the ground for recommendations.
@@ 4.1 Approach
The construction of our benchmark dataset, illustrated in Fig. 2, centers on three key components: NL profiles, candidate items, and ground truth. Our methodology begins by sampling authors from a scholarly dataset (arXiv). For each author, their published papers are temporally divided into an earlier and a later portion. The earlier portion is utilized for generating an NL interest profile through LLM prompting, while the later portion is utilized for creating the ground truth set of recommendations. For each author, we also sample additional papers from arXiv to create a set of candidate items. The subsequent recommendation task involves ranking these candidate papers with respect to the NL profile. Further details on each step are provided below.
@@ 4.1.1 Scholarly Dataset
We use the public arXiv dataset released by Clement et al. [8]. [^1] https://github.com/mattbierbaum/arxiv-public-datasets/releases/tag/v0.2.0 It consists of article metadata for 1,506,500 articles, including the arXiv paper ID, author list, title, abstract, version, and category listings. Additionally, it provides a citation network of papers. The dataset covers articles published up to March 1, 2019. We performed author name normalization, assigned unique author IDs, and retained only authors with a valid name who had published at least 10 papers. We only kept articles that had metadata and were also present in the citation network.
@@ 4.1.2 NL Profile Generation
We sample n=1,000n{=}1,000 authors from the filtered arXiv dataset. For each author $u$, let $\mathcal{I}_{u}$ denote the articles they authored. We temporarily order $\mathcal{I}_{u}$ and divide to earlier portion, $\mathcal{I}_{u}^{\blacktriangle}$ (for NL profile construction), and later portion, $\mathcal{I}_{u}^{\blacktriangledown}$ (for ground truth creation). For simplicity, we divide $\mathcal{I}_{u}$ in half to ensure that there are enough articles for both NL profile generation and ground truth creation; exploring other division strategies is left for future work.
@@ 4.1.3 Ground Truth
The later portion of the articles, $\mathcal{I}_{u}^{\blacktriangledown}$, which were not used for NL profile generation, are utilized for ground truth creation. We consider all papers referenced by any article in $\mathcal{I}_{u}^{\blacktriangledown}$ as ground truth items, provided they were not “seen” by the author before, that is, referenced by any of the papers in $\mathcal{I}_{u}^{\blacktriangle}$ or authored by the individual themselves (i.e., exclude self-citations). We denote this set of ground truth items as $\mathcal{G}_{u}$. In this initial approach, we do not distinguish between the varying degrees of relevance among these items.
@@ 4.2 Characterizing Profile Breadth
#Candidate items per author 1,000
#Ground truth papers per author (min/median/max) 1/27/438
@@ 4.3 Dataset Statistics
The key statistics for the SciNUP dataset are summarized in Table 2. With 1,000 authors, the dataset supports robust statistical analysis, and the diversity in paper counts offers a realistic and varied set of author profiles. The most frequently represented arXiv categories in which the authors in our dataset publish are diverse fields of physics, mathematics, and computer science; see Figure 5(a).
@@ 7 Conclusion
Our work also opens several exciting avenues for the task of NL profile generation. The construction process involves various parameters and prompting strategies that can be further explored, such as the ratio of papers used for profile creation versus evaluation and and the specific instructions given to the LLM. Furthermore, validating the generated NL profiles through human evaluation would provide crucial insights into their quality and fidelity to actual user interests, guiding the development of more effective profile generation techniques.

### v2

@@ 4 The SciNUP Dataset
We present SciNUP (Scientific Natural Language User Profiles), a synthetic dataset designed for evaluating natural language-based recommendation in the scholarly domain. We simulate a real-world scenario in which a researcher—hereafter referred to as the user—is looking for article recommendations relevant to their specific expertise and interests. To model this authentically, we leverage the user’s publication history. We utilize their past authored papers to create NL user interest profiles; subsequently, the cited references in their later publications serve as the ground truth for evaluating recommendation effectiveness.
@@ 4.1 Approach
The construction of our benchmark dataset, illustrated in Fig. 2, centers on three key components: NL profiles, candidate items, and ground truth. Our methodology begins by sampling users (authors) from a scholarly dataset (arXiv). For each user, their published papers are temporally divided into an earlier and a later portion. The earlier portion is utilized for generating an NL interest profile through LLM prompting, while the later portion is utilized for creating the ground truth set of recommendations. For each user, we also sample additional papers from arXiv to create a set of candidate items. The subsequent recommendation task involves ranking these candidate papers with respect to the NL profile. Further details on each step are provided below.
@@ 4.1.1 Scholarly Dataset
We use the public arXiv dataset released by Clement et al. [8]. [^1] https://github.com/mattbierbaum/arxiv-public-datasets/releases/tag/v0.2.0 To ensure maximum accessibility and reproducibility, we rely exclusively on article metadata, which is released under a CC0 license. This metadata includes the arXiv paper ID, author list, title, abstract, version, and category listings for 1,506,500 articles. Furthermore, we integrate a citation network provided by the same source to establish relational links between papers. The dataset covers articles published up to March 1, 2019. We performed author name normalization, assigned unique user IDs, and retained only users with a valid name who had published at least 10 papers. We only kept articles that had metadata and were also present in the citation network.
@@ 4.1.2 NL Profile Generation
We sample n=1,000n{=}1,000 users from the filtered arXiv dataset. For each user $u$, let $\mathcal{I}_{u}$ denote the articles they authored. We temporarily order $\mathcal{I}_{u}$ and divide to earlier portion, $\mathcal{I}_{u}^{\blacktriangle}$ (for NL profile construction), and later portion, $\mathcal{I}_{u}^{\blacktriangledown}$ (for ground truth creation). For simplicity, we divide $\mathcal{I}_{u}$ in half to ensure that there are enough articles for both NL profile generation and ground truth creation; exploring other division strategies is left for future work.
@@ 4.1.3 Ground Truth
The later portion of the articles, $\mathcal{I}_{u}^{\blacktriangledown}$, which were not used for NL profile generation, are utilized for ground truth creation. We consider all papers referenced by any article in $\mathcal{I}_{u}^{\blacktriangledown}$ as ground truth items, provided they were not “seen” by the user before, that is, referenced by any of the papers in $\mathcal{I}_{u}^{\blacktriangle}$ or authored by the individual themselves (i.e., exclude self-citations). We denote this set of ground truth items as $\mathcal{G}_{u}$. In this initial approach, we do not distinguish between the varying degrees of relevance among these items.
@@ 4.2 Characterizing Profile Breadth
#Candidate items per user 1,000
#Ground truth papers per user (min/median/max) 1/27/438
@@ 4.3 Dataset Statistics
The key statistics for the SciNUP dataset are summarized in Table 2. With 1,000 users, the dataset supports robust statistical analysis, and the diversity in paper counts offers a realistic and varied set of user profiles. The most frequently represented arXiv categories in which the users in our dataset publish are diverse fields of physics, mathematics, and computer science; see Figure 5(a).
@@ 7 Conclusion
Our work also opens several exciting avenues for the task of NL profile generation. The construction process involves various parameters and prompting strategies that can be further explored, such as the ratio of papers used for profile creation versus evaluation and the specific instructions given to the LLM. Furthermore, validating the generated NL profiles through human evaluation would provide crucial insights into their quality and fidelity to actual user interests, guiding the development of more effective profile generation techniques.
7.0.1 Acknowledgements
This research was supported by the Norwegian Research Center for AI Innovation, NorwAI (Research Council of Norway, project number 309834).
7.0.2 Disclosure of Interests.
The authors have no competing interests to declare that are relevant to the content of this article.